# Vertex-based Graph Neural Solver and its Application to Linear Elasticity Equations

### v1

@@ 4.2 Method 2: Adaptive Graph Fourier Neural Solver (AG-FNS)
To overcome the limitations of G-FNS in handling variable-coefficient problems, we enhanced its frequency domain correction operator $\mathcal{H}$ by introducing a learnable network $\mathcal{M}$. This network learns adaptive Fourier basis functions to accommodate the multi-scale characteristics introduced by variable coefficients. The improved $\mathcal{H}$ is defined as:
$\mathcal{H}=\mathcal{F}^{-1}\bigl(\boldsymbol{\xi}(\boldsymbol{x})\bigr)\,\mathcal{C}^{*}\,\tilde{\Lambda}\,\mathcal{C}\,\mathcal{F}\bigl(\boldsymbol{\xi}(\boldsymbol{x})\bigr),$ (25)
where $\boldsymbol{\xi}(\boldsymbol{x})$ is a mapping of the original coordinates $\boldsymbol{x}$ via the network $\mathcal{M}$:
$\boldsymbol{\xi}(\boldsymbol{x})=\mathcal{M}(\boldsymbol{x})=\boldsymbol{x}+\boldsymbol{x}\odot f_{\theta}(\boldsymbol{x}),$ (26)
where $f_{\theta}$ is a learnable fully connected network (see [45] for the architecture).
The mapping $\mathcal{M}$ can be interpreted as a learned basis transformation. Unlike the traditional Fourier transform, which relies on a fixed basis $e^{i\mathbf{k}\cdot\mathbf{x}_{l}}$, the transform incorporating $\mathcal{M}$ employs an adaptive basis $e^{i\mathbf{k}\cdot\boldsymbol{\xi}_{l}}$:
$\hat{\boldsymbol{r}}^{i}(\mathbf{k})=\sum_{l=1}^{N}\boldsymbol{r}_{l}^{i}e^{i\mathbf{k}\cdot\boldsymbol{\xi}_{l}},\quad\mathbf{k}\in[-m,\dots,m]^{d},$ (27)
@@
Figure 8: Schematic diagram of the adaptive operator $\mathcal{H}$ in AG-FNS. The key enhancement over the G-FNS is the introduction of the coordinate transformation path (top panel): physical coordinates $\boldsymbol{x}$ are mapped by the learnable network $\mathcal{M}$ to latent coordinates $\boldsymbol{\xi}$. These adaptive coordinates $\boldsymbol{\xi}$ are explicitly injected into the Fourier transform ($\mathcal{F}$) and inverse transform ($\mathcal{F}^{-1}$) blocks to construct dynamic, problem-dependent basis functions, allowing the solver to accommodate variable coefficients effectively.
@@
We further explore the performance of AG-FNS in a 3D scenario (Data-2). The number of nodes is $N=3685$ (11055 DOFs). Due to the weaker smoothing effect in 3D problems, we increase the operator $\mathcal{B}$ to $50$ iterations of the block weighted Jacobi method. The frequency parameter is $m=6$. The blue solid line in Figure 11(a) shows the training loss, which decreases slowly. Test results (the purple line in Figure 11(b)(c)) indicate that AG-FNS cannot achieve the desired convergence speed in this 3D case, suggesting that the single-level approach struggles to effectively handle complex high-dimensional problems and requires further refinement.
@@ 4.3 Method 3: Multilevel Adaptive Graph Fourier Neural Solver (ML-AG-FNS)
This method constructs a hierarchy of correction operators $\mathcal{H}=\{\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{L}\}$. The definition of the $i$-th level operator $\mathcal{H}_{i}$ is similar to Equation (25), but employs different frequency domain parameters $m_{i}$, satisfying the hierarchy $m_{L}<\cdots<m_{1}$.
$\mathcal{H}_{i}=\mathcal{F}_{i}^{-1}\bigl(\boldsymbol{\xi}(\boldsymbol{x})\bigr)\,\mathcal{C}_{i}^{*}\,\tilde{\Lambda}_{i}\,\mathcal{C}_{i}\,\mathcal{F}_{i}\bigl(\boldsymbol{\xi}(\boldsymbol{x})\bigr),\qquad i=1,\ldots,L.$
@@
Figure 12: Schematic diagram of the ML-AG-FNS workflow. The iteration begins with pre-smoothing the current solution $\boldsymbol{u}^{k}$ to obtain $\tilde{\boldsymbol{u}}^{k}$. This is followed by a cascade of $L$ correction levels, where each level $i$ computes a residual $\boldsymbol{r}_{i}^{k}$ and applies a specific operator block $\mathcal{H}_{i}+\mathcal{B}$ (configured with distinct frequency parameters $m_{i}$) to generate the correction term $\boldsymbol{e}_{i}^{k}$. The arrows between levels indicate that the residual for the subsequent level is updated based on the previous corrections. Finally, all hierarchical corrections $\boldsymbol{e}_{1}^{k},\dots,\boldsymbol{e}_{L}^{k}$ are aggregated and added to the base iterate to produce the final updated solution $\boldsymbol{u}^{k+1}$.
@@
In the Data-3 case, we employ $50$ block weighted Jacobi iterations for $\mathcal{B}$, with $L=4$ and frequency parameters $m_{i}=21-i$. Figure 13 shows the residual reduction curve on a single test sample. It is evident that while traditional SA-AMG exhibits limited convergence speed when solving these anisotropic problems, ML-AG-FNS maintains a rapid convergence rate. This indicates that the method can adaptively capture the complex error propagation characteristics induced by material rotation and anisotropy.
@@
5 Conclusion and Discussion
This paper proposes and develops the ML-AG-FNS, a novel solver designed to tackle large-scale sparse linear algebraic systems arising from the discretization of variable-coefficient PDE systems on unstructured grids, with the linear elasticity model serving as a prime example. This work successfully addresses the inherent limitations of FNS, specifically overcoming their dependency on structured grids and vector-valued problems. Numerical experiments conducted on a diverse suite of complex linear elasticity datasets—encompassing 2D and 3D domains, isotropic and strongly anisotropic conditions, as well as heterogeneous materials—demonstrate the superior performance of the proposed method. Whether deployed as a standalone solver or as a preconditioner, ML-AG-FNS exhibits convergence rates and robustness that significantly surpass those of optimized traditional SA-AMG methods. Notably, this performance advantage is particularly pronounced when addressing challenging anisotropic problems.
In the future, we plan to conduct further research in the following areas:
  1. 1.
    Theoretical Analysis and Interpretability: While ML-AG-FNS has empirically demonstrated robust convergence and efficiency, establishing a rigorous theoretical framework remains essential. Future work will aim to analyze the convergence mechanism in depth, for instance, by attempting to extend Local Fourier Analysis (LFA) to the context of linear elasticity equations. Simultaneously, we will explore the relationship between the effective adaptive basis learned by the mapping $\mathcal{M}$ and the spectral properties of the discrete operator, providing a stronger theoretical foundation for the method’s effectiveness.
  2. 2.
    Optimization of Training Strategies: The current training objective, which focuses on the convergence of the fixed-point iteration, has proven sufficient to yield a highly effective solver. To further unlock the method’s potential, future work will investigate end-to-end training strategies that directly optimize the convergence performance of the preconditioned FGMRES. We anticipate that this direct optimization could lead to even superior acceleration and stability.
  3. 3.
    Computational Efficiency and Large-Scale Scalability: The proposed GNN architecture has successfully managed the complexity of unstructured topologies on moderate-scale problems. Building on this foundation, we plan to further optimize the implementation efficiency of the GNN architecture and message-passing mechanisms. We will also evaluate the performance of ML-AG-FNS on large-scale practical engineering problems (e.g., exceeding ten million degrees of freedom) and investigate its scalability in distributed parallel computing environments.
  4. 4.
    Generalization to More Complex Physical Systems: The ML-AG-FNS framework has shown exceptional adaptability to vector-valued problems with complex coefficients. Leveraging this versatility, we aim to extend the proposed multilevel adaptive FNS framework to other types of complex PDE systems, including nonlinear elastoplastic mechanics, electromagnetics (Maxwell’s equations), and multi-physics coupling problems such as fluid-structure interaction.
@@ Acknowledgments
This work is funded by the NSFC grants (12371373). Shi Shu is supported by Science Challenge Project (TZ2024009). This work was carried out in part using computing resources at the High Performance Computing Platform of Xiangtan University.

### v2

@@ 5.2 Method 2: Adaptive Graph Fourier Neural Solver (AG-FNS)
To overcome the limitations of G-FNS in handling variable-coefficient problems, we enhanced its frequency domain correction operator $\mathcal{H}$ by introducing a learnable network $\mathcal{M}$. This network learns adaptive Fourier basis functions to accommodate the multi-scale characteristics introduced by variable coefficients. The improved $\mathcal{H}$ is defined as
$\mathcal{H}=\mathcal{F}^{-1}\bigl(\bm{\xi}(\bm{x})\bigr)\,\mathcal{C}^{*}\,\tilde{\Lambda}\,\mathcal{C}\,\mathcal{F}\bigl(\bm{\xi}(\bm{x})\bigr),$ (42)
where $\bm{\xi}(\bm{x})$ is a mapping of the original coordinates $\bm{x}$ via the network $\mathcal{M}$
$\bm{\xi}(\bm{x})=\mathcal{M}(\bm{x})=\bm{x}+\bm{x}\odot f_{\theta}(\bm{x}),$ (43)
where $f_{\theta}$ is a learnable fully connected network (see [li2023fourier] for the architecture).
The mapping $\mathcal{M}$ can be interpreted as a learned basis transformation. Unlike the traditional Fourier transform, which relies on a fixed basis $e^{i\bm{k}\cdot\bm{x}_{l}}$, the transform incorporating $\mathcal{M}$ employs an adaptive basis $e^{i\bm{k}\cdot\bm{\xi}_{l}}$:
$\hat{\bm{r}}^{(i)}(\bm{k})=\sum_{l=1}^{N}\bm{r}_{l}^{(i)}e^{i\bm{k}\cdot\bm{\xi}_{l}},\quad\bm{k}\in[-m,\dots,m]^{d},$ (44)
@@
Figure 8: Schematic diagram of the adaptive operator $\mathcal{H}$ in AG-FNS. The key enhancement over the G-FNS is the introduction of the coordinate transformation path (top panel): physical coordinates $\bm{x}$ are mapped by the learnable network $\mathcal{M}$ to latent coordinates $\bm{\xi}$. These adaptive coordinates $\bm{\xi}$ are explicitly injected into the Fourier transform ($\mathcal{F}$) and inverse transform ($\mathcal{F}^{-1}$) blocks to construct dynamic, problem-dependent basis functions, allowing the solver to accommodate variable coefficients effectively.
@@
We further explore the performance of AG-FNS in a 3D scenario (Data-2). The number of nodes is $N=3685$ (11055 DOFs). Due to the weaker smoothing effect in 3D problems, we increase the operator $\mathcal{B}$ to $50$ iterations of the weighted block Jacobi Method. The frequency parameter is $m=6$. The blue solid line in Figure 11(a) shows the training loss, which decreases slowly. Test results (the purple line in Figure 11(b)(c)) indicate that AG-FNS cannot achieve the desired convergence speed in this 3D case, suggesting that the single-level approach struggles to effectively handle complex high-dimensional problems and requires further refinement.
@@ 5.3 Method 3: Multilevel Adaptive Graph Fourier Neural Solver (ML-AG-FNS)
This method constructs a hierarchy of correction operators $\mathcal{H}=\{\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{L}\}$. The definition of the $i$-th level operator $\mathcal{H}_{i}$ is similar to (42), but employs different frequency domain parameters $m_{i}$, satisfying the hierarchy $m_{L}<\cdots<m_{1}$.
$\mathcal{H}_{i}=\mathcal{F}_{i}^{-1}\bigl(\bm{\xi}(\bm{x})\bigr)\,\mathcal{C}_{i}^{*}\,\tilde{\Lambda}_{i}\,\mathcal{C}_{i}\,\mathcal{F}_{i}\bigl(\bm{\xi}(\bm{x})\bigr),\qquad i=1,\ldots,L.$
@@
Figure 12: Schematic diagram of the ML-AG-FNS workflow. The iteration begins with pre-smoothing the current solution $\bm{u}^{(k)}$ to obtain $\bm{u}^{(k+\frac{1}{2})}$. This is followed by a cascade of $L$ correction levels, where each level $i$ computes a residual $\bm{r}_{i}^{(k+\frac{1}{2})}$ and applies a specific operator block $\mathcal{H}_{i}+\mathcal{B}$ (configured with distinct frequency parameters $m_{i}$) to generate the correction term $\bm{e}_{i}^{(k+\frac{1}{2})}$. The arrows between levels indicate that the residual for the subsequent level is updated based on the previous corrections. Finally, all hierarchical corrections $\bm{e}_{1}^{(k+\frac{1}{2})},\dots,\bm{e}_{L}^{(k+\frac{1}{2})}$ are aggregated and added to $\bm{u}^{(k+\frac{1}{2})}$ to produce the final updated solution $\bm{u}^{(k+1)}$.
@@
In the Data-3 case, we employ $50$ weighted block Jacobi iterations for $\mathcal{B}$, with $L=4$ and frequency parameters $m_{i}=21-i$. Figure 13 shows the residual reduction curve on a single test sample. It is evident that while traditional SA-AMG exhibits limited convergence speed when solving these anisotropic problems, ML-AG-FNS maintains a rapid convergence rate. This indicates that the method can adaptively capture the complex error propagation characteristics induced by material rotation and anisotropy.
@@
6 Conclusion and Future Work
In this paper, we present a deep learning-based hybrid iterative framework designed to solve variable-coefficient linear elasticity equations on arbitrary unstructured meshes. By bridging the gap between classical numerical linear algebra and geometric deep learning, we addressed the limitations of traditional smoothers in handling the complex spectral characteristics of vector-valued PDE systems. Our methodology evolved from a basic graph-based solver to the ML-AG-FNS. The core innovation lies in the integration of a learnable coordinate transformation, which constructs a dynamic, problem-dependent spectral basis capable of capturing multi-scale error modes induced by material heterogeneity. Furthermore, the introduction of a multilevel hierarchy enables the solver to perform effective error correction across cascading frequency bandwidths, thereby overcoming the scalability issues encountered in high-dimensional and strongly anisotropic problems. Theoretical analysis, grounded in Korn’s inequality, provides a rigorous guarantee of unconditional convergence for the proposed hybrid scheme. Extensive numerical experiments confirm that our method significantly outperforms the SA-AMG method in terms of convergence rate and robustness, particularly in challenging 3D orthotropic and general anisotropic scenarios.
Looking forward, several promising directions for future research emerge:
  • Extension to Nonlinear Mechanics: We plan to extend the current framework to handle nonlinear elasticity problems, such as hyperelasticity and plasticity. The hybrid iterative structure offers a potential pathway to accelerate Newton-Raphson iterations by effectively preconditioning the linearized tangent stiffness matrix at each step.
  • Transfer Learning and Generalization: While the current model trains effectively on specific distributions of material parameters, enhancing its zero-shot generalization capabilities across entirely different geometric topologies or boundary conditions via meta-learning or transfer learning remains a key objective.
  • Large-Scale Parallelization: To tackle industrial-scale engineering problems with millions of degrees of freedom, we aim to investigate the parallel implementation of the proposed graph neural operators on distributed computing clusters, optimizing the trade-off between communication costs and inference speed.
  • Multi-Physics Coupling: We intend to generalize the vector-valued solver architecture to other coupled systems, such as fluid-structure interaction or poroelasticity, where the block-coupled nature of the governing equations presents similar numerical challenges.
@@ Acknowledgments
This work is funded by the NSFC grants (12371373). Shi Shu is supported by the Science Challenge Project (TZ2024009). Yun Liu is supported by the Postgraduate Scientific Research Innovation Project of Xiangtan University (XDCX2025Y193). Zhen Wang is supported by the Hunan Provincial Innovation Foundation for Postgraduates (CX20230616). Computations were performed at the High Performance Computing Platform of Xiangtan University.